# Re-accretion of Giant Impact Ejecta Can Drive Significant Atmospheric Erosion on Terrestrial Planets

### v1

@@ 1 Introduction
Atmospheric growth can occur through processes such as mantle degassing and the delivery of volatiles by comets and asteroids, both of which are thought to have contributed to the formation of Earth’s current atmosphere (e.g., Elkins-Tanton, 2008; Lebrun et al., 2013; Avice et al., 2017; Wyatt et al., 2020; Sinclair et al., 2020). Conversely, atmospheric erosion, driven by processes including photo-evaporation and/or core-powered mass loss, plays a critical role in sculpting the observed radius distribution of short-period exoplanets (e.g., Owen & Wu, 2017; Ginzburg et al., 2018). Giant impacts during the final stage of planetary accretion are also a driver of atmospheric loss (e.g., Schlichting et al., 2015; Schlichting & Mukhopadhyay, 2018; Kegerreis et al., 2020a; Izidoro et al., 2022), but their efficiency has been historically underestimated as past studies have primarily considered their immediate effects, neglecting longer-term consequences. In this study, we revisit the efficiency of impact-driven atmospheric loss following giant impacts.
@@
In the final phase of planet formation, Mars-sized planetary embryos grow via colliding among themselves in giant impacts, driven by a chaotic phase of dynamical instabilities, which shapes the orbital and structural properties of the planets we observe today (e.g., Gabriel & Cambioni, 2023). In our solar system, such impacts are hypothesized to have led to the formation of Earth’s Moon (e.g., Hartmann & Davis, 1975; Benz et al., 1986; Canup & Asphaug, 2001; Canup, 2012; Ćuk & Stewart, 2012) and stripped the mantle of Mercury (e.g., Benz et al., 1988, 2007; Chau et al., 2018). In extrasolar systems, it now well-established that instabilities play a key role in shaping the orbital architectures (e.g., Rasio & Ford, 1996; Chatterjee et al., 2008; Izidoro et al., 2017; Frelikh et al., 2019; Ghosh & Chatterjee, 2023; Chen et al., 2025), with giant impacts contributing to the ‘peas-in-a-pod’ trend observed in the Kepler systems (Ghosh & Chatterjee, 2023; Lammers et al., 2023). Giant impacts are also thought to be responsible for creating the high bulk density planets known as super-Mercuries (e.g., Reinhardt et al., 2022; Dou et al., 2024) as well as the density disparities, observed in some of the adjacent planet pairs (e.g., Liu et al., 2015; Inamdar & Schlichting, 2016; Bonomo et al., 2019).
@@
Most giant impacts between protoplanets do not result in perfect mergers but rather lead to partial accretion and hit-and-run encounters (e.g., Agnor & Asphaug, 2004; Asphaug et al., 2006; Genda et al., 2012; Leinhardt & Stewart, 2012). The outcomes of these collisions have been extensively studied through numerical simulations in various contexts, ranging from the canonical Moon-forming impact (e.g., Benz et al., 1986; Canup & Asphaug, 2001; Canup et al., 2013; Kegerreis et al., 2022) to the collisions between newly discovered super-Earth and sub-Neptune class planets (e.g., Marcus et al., 2009, 2010; Liu et al., 2015; Hwang et al., 2017, 2018; Kegerreis et al., 2018, 2020a; Denman et al., 2020, 2022; Ghosh et al., 2024). Such systematic investigations of giant impacts across a broad parameter space, including variations in impact angle, impact velocity, and the mass and composition of both the impactor and the target, provide valuable insights into the violent final stage of planetary accretion. These simulations constrain the amount of mass ejected during the impact (e.g., Leinhardt & Stewart, 2012; Cambioni et al., 2019; Emsenhuber et al., 2020) and have also led to the development of scaling laws to estimate the extent of atmospheric erosion in such events (e.g., Shuvalov, 2009; Kegerreis et al., 2020b; Denman et al., 2022). These studies commonly suggest that only partial atmospheric erosion occurs in typical giant impacts. For instance, the atmospheric mass loss prescription from Kegerreis et al. (2020b) estimates that only $10\%$ of Earth’s primordial atmosphere could have been lost from the immediate effects of the canonical Moon forming impact.
@@
Due to the prohibitive computational cost associated with these numerical simulations, the primary focus of these studies has typically been on the impact and its immediate aftermath, typically covering only a few days. However, these impacts eject a substantial amount of mass from the colliding bodies, some of which persists within the system long after the primary impact (e.g., Jackson & Wyatt, 2012; Jackson et al., 2014; Genda et al., 2015; Watt et al., 2024). Depending on the specifics of the impact, some fraction of this ejected mass would be vaporised (e.g., Watt et al., 2021; Nakajima et al., 2022). As the vaporised ejecta expands, it cools and condenses, producing dust at detectable levels, which are then rapidly depleted to non-detectable levels through collisional evolution and subsequently removed from the system via radiation forces (e.g., Wyatt & Jackson, 2016; Watt et al., 2021). The non-vaporised debris survives much longer and will gradually erode through mutual collisions and other physical processes, including gravitational perturbations from the surviving planet(s) (e.g., Jackson & Wyatt, 2012; Watt et al., 2024). Over time, these objects will be progressively ground down to micrometre-sized dust through collisional evolution and eventually removed by radiation pressure or Poynting-Robertson drag (e.g., Wyatt, 2008; Wyatt et al., 2011). This theoretical understanding is supported by observations of extrasolar systems with large quantities of warm dust in the terrestrial region, inferred to have been recently created in collision between the planetary embryos (e.g., Lisse et al., 2009; Weinberger et al., 2011; Genda et al., 2015; Moór et al., 2021; Schneiderman et al., 2021; Watt et al., 2024). Importantly, during this long process, a fraction of the debris will re-impact the planet and be re-accreted into the remaining planet(s) (Jackson & Wyatt, 2012; Wyatt et al., 2017), potentially causing further atmospheric erosion.
@@ 2 Methods
In this study, we adopt a simplified approach to investigate the long-term consequences of giant impacts. Depending on the kinematics, giant impacts can result in a range of different outcomes, from efficient accretion and hit-and-run collisions to catastrophic disruptions (e.g., Leinhardt & Stewart, 2012; Emsenhuber et al., 2020). But regardless of the outcome, a substantial amount of unbound mass is ejected during these impacts. This unbound mass is generally a mixture of vapour and solid material [^1] By solid material, we refer to all the non-vaporised condensed matter, which may eventually cool to become solid debris. with varying proportions, depending on the specifics of the impact (e.g., Ćuk & Stewart, 2012; Carter et al., 2020; Nakajima et al., 2022; Watt et al., 2021). However, in this study, our focus is not on the detailed outcomes of individual giant impact events. Instead, we attempt to parametrize the outcome in a simplified manner that is broadly applicable across a range of giant impacts. Hence, we adopt a simplified model in which a giant impact results in the ejection of a fraction $f_{\rm{esc}}$} of the surviving planet’s mass $m_{\rm{p}}$} as unbound solid material and a fraction $f_{\rm{esc,v}}$} as vapor. The unbound ejecta would expand out from the point of collision and shear out to eventually form an axisymmetric debris disk around the host star encompassing the planet and the collision point (e.g., Jackson & Wyatt, 2012; Jackson et al., 2014; Genda et al., 2015; Watt et al., 2024). The debris are assumed to have some velocity dispersion in addition to the Keplerian velocity of the planet.
@@ 2.1 Evolution of Debris from Giant Impacts & Re-accretion
where, $M_{\rm{*}}$} is the stellar mass, $a_{\rm{p}}$} is the semi-major axis, and $\rho_{\rm{p}}$} is the bulk density of the planet. Throughout the study, we have assumed that the debris has a density of ${\rho_{\rm{imp}}}=2.7\ {{\rm g\ cm^{-3}}}$, and a dispersal threshold ${Q_{\rm{D}}^{*}}=10^{5}\ {{\rm J\ kg^{-1}}}$. The size of the largest objects in the debris ($D_{\rm{max}}$}) from a giant impact is highly uncertain. The vaporised component of the ejecta would expand and cool to form condensates with sizes ($D_{\rm{vc}}$}) in the millimetre to centimetre range (e.g., Melosh & Vickery, 1991; Johnson & Melosh, 2014), while the non-vaporised material may reach up to several hundred kilometres. For instance, Genda et al. (2015) found that the largest debris fragments [^3] Here, debris fragments refer to ejected material and should not be confused with the remnant post-impact protoplanet(s). could be as massive as half the lunar mass, corresponding to a diameter of $\sim 1380\ {\rm{km}}$. More recently, Watt et al. (2024) found that the debris fragments are generally larger than $\sim 100\ {\rm{km}}$ in diameter. This suggests that the size distribution of ejecta should consist of two primary components, the smaller vapour condensates and the larger non-vaporised material. Given the substantial differences in their sizes, we expect each component to set up a collisional cascade and to evolve essentially independently of the other. The small vapour condensates would have short collisional lifetimes, and are therefore rapidly depleted from the system, typically within a small number of orbits (see Figure 5 of Wyatt & Jackson, 2016). Due to their small size, these condensates are ineffective in atmospheric erosion (e.g., Schlichting et al., 2015) and thus would have a negligible influence on the post-impact planet(s) over their short lifetimes. In contrast, the larger non-vaporised debris would have much longer collisional lifetimes (e.g., $\tau_{c}\sim 6$ Myr for ${D_{\rm{max}}}=100\ {\rm{km}}$, considering an Earth-like progenitor at $1\ {\rm{au}}$ and ${M_{\rm{d}}}=0.01\ {M_{\earth}}$), increasing the likelihood of re-accretion and potentially influencing the planet’s subsequent evolution. Therefore, we exclude the vaporised component from our analysis and, for the remainder of this study, focus exclusively on the non-vaporised debris, adopting ${M_{\rm{d}}}={f_{\rm{esc}}}{m_{\rm{p}}}$.
@@ 3 Results
We first consider the possible implications of debris re-accretion on Earth following a giant impact such as that hypothesised to have formed the Moon. This extensively studied event, observationally constrained by the Earth-Moon configuration (e.g., Canup & Asphaug, 2001; Timpe et al., 2023; Zhou et al., 2024), is a natural reference point for our study. We note, however, that the exact nature of the impact scenario that led to the formation of the Moon remains uncertain and debated. It is not the purpose of our study to address these, but to adopt one representative scenario as a reference. While we use the giant impact hypothesis of the Moon formation as a reference point, the focus of our work is not restricted to satellite-forming impacts. Instead, we aim to explore the broader role of typically overlooked debris in causing atmospheric erosion through secondary impacts, regardless of whether the primary impact forms a satellite or not. To this end, we subsequently expand our analysis to explore a broader range of parameter space.
@@ 3.1.3 Effect of Initial Atmospheric Properties
In the previous sections, we assumed that Earth, after the giant impact, had an atmosphere similar to that of the present-day Earth. In this section, we explore the long-term atmospheric erosion following the impact, focusing on the effects of initial atmospheric mass and mean molecular weight on this process.
@@
The mean molecular weight of the atmosphere, $\mu$, directly affects the atmospheric scale height $H$, while the atmospheric mass primarily influences the base density, $\rho_{\rm{b}}$}. For a given planet and impact velocity, the values of $H$ and $\rho_{\rm{b}}$} determine the size range of the impactors that most efficiently erode the atmosphere. In the atmospheric erosion prescription of Shuvalov (2009), this relationship is encapsulated in the definition of the erosional efficiency factor, $\xi$ (section 2.2), which determines the extent of atmospheric erosion (Equation 11). For a given impact velocity and impactor mass, atmospheric mass loss peaks at $\xi\sim 350$ and falls for both higher and lower values of $\xi$, declining to less than $10\%$ of its peak value beyond the range $10\lesssim\xi\lesssim 10^{5}$ (see Figure 6 of Shuvalov (2009)). From section 2.2, we find that $D^{3}\propto\xi{m_{\rm{atm}}}/\mu^{2}$. So, as $m_{\rm{atm}}$} increases, the efficient erosion regime ($10\lesssim\xi\lesssim 10^{5}$) shifts towards the larger impactor sizes, where we have more available mass in the debris, based on the adopted debris size distribution (section 2.1). If the efficient erosion regime remains below the maximum debris size ($D_{\rm{max}}$}), the atmospheric erosion per unit accreted mass increases (Equation 14). However, a continued increase in $m_{\rm{atm}}$} may shift the efficient erosion regime to even larger impactor sizes exceeding $D_{\rm{max}}$}, leading to a decline in atmospheric erosion. This transition occurs at lower $m_{\rm{atm}}$} for lower values of $\mu$ (e.g., at ${m_{\rm{atm}}}\simeq 0.7\ {m_{\rm{atm\earth}}}$ for $\mu=2.35$, while at ${m_{\rm{atm}}}\simeq 100\ {m_{\rm{atm\earth}}}$ for $\mu=29$) due to the $\mu$ dependence of $\xi$.

### v2

@@ 1 Introduction
Atmospheric growth can occur through processes such as mantle degassing and the delivery of volatiles by comets and asteroids, both of which are thought to have contributed to the formation of Earth’s current atmosphere (e.g., Elkins-Tanton, 2008; Lebrun et al., 2013; Avice et al., 2017; Wyatt et al., 2020; Sinclair et al., 2020). Conversely, atmospheric erosion, driven by processes including photo-evaporation and/or core-powered mass loss, plays a critical role in sculpting the observed radius distribution of short-period exoplanets (e.g., Owen & Wu, 2017; Ginzburg et al., 2018). Giant impacts during the final stage of planetary accretion are also a driver of atmospheric loss (e.g., Genda & Abe, 2003, 2005; Schlichting et al., 2015; Schlichting & Mukhopadhyay, 2018; Kegerreis et al., 2020a, b; Denman et al., 2020, 2022; Izidoro et al., 2022; Ghosh et al., 2024; Lock & Stewart, 2024; Roche et al., 2025), but their efficiency has been historically underestimated as past studies have primarily considered their immediate effects, neglecting longer-term consequences. In this study, we revisit the efficiency of impact-driven atmospheric loss following giant impacts.
@@
In the final phase of planet formation, Mars-sized planetary embryos grow via colliding among themselves in giant impacts, driven by a chaotic phase of dynamical instabilities, which shapes the orbital and structural properties of the planets we observe today (e.g., Gabriel & Cambioni, 2023). In our solar system, such impacts are hypothesized to have led to the formation of Earth’s Moon (e.g., Cameron & Ward, 1976; Hartmann & Davis, 1975; Benz et al., 1986; Canup & Asphaug, 2001; Canup, 2012; Ćuk & Stewart, 2012; Lock et al., 2018) and stripped the mantle of Mercury (e.g., Benz et al., 1988, 2007; Chau et al., 2018). In extrasolar systems, it now well-established that instabilities play a key role in shaping the orbital architectures (e.g., Rasio & Ford, 1996; Chatterjee et al., 2008; Izidoro et al., 2017; Frelikh et al., 2019; Ghosh & Chatterjee, 2023; Chen et al., 2025), with giant impacts contributing to the ‘peas-in-a-pod’ trend observed in the Kepler systems (Ghosh & Chatterjee, 2023; Lammers et al., 2023). Giant impacts are also thought to be responsible for creating the high bulk density planets known as super-Mercuries (e.g., Reinhardt et al., 2022; Dou et al., 2024) as well as the density disparities, observed in some of the adjacent planet pairs (e.g., Liu et al., 2015; Inamdar & Schlichting, 2016; Bonomo et al., 2019).
@@
Most giant impacts between protoplanets do not result in perfect mergers but rather lead to partial accretion and hit-and-run encounters (e.g., Agnor & Asphaug, 2004; Asphaug et al., 2006; Genda et al., 2012; Leinhardt & Stewart, 2012). The outcomes of these collisions have been extensively studied through numerical simulations in various contexts, ranging from the canonical Moon-forming impact (e.g., Benz et al., 1986; Canup & Asphaug, 2001; Canup et al., 2013; Kegerreis et al., 2022) to the collisions between newly discovered super-Earth and sub-Neptune class planets (e.g., Marcus et al., 2009, 2010; Liu et al., 2015; Hwang et al., 2017, 2018; Denman et al., 2020, 2022; Ghosh et al., 2024; Roche et al., 2025). Such systematic investigations of giant impacts across a broad parameter space, including variations in impact angle, impact velocity, and the mass and composition of both the impactor and the target, provide valuable insights into the violent final stage of planetary accretion. These simulations constrain the amount of mass ejected during the impact (e.g., Leinhardt & Stewart, 2012; Cambioni et al., 2019; Emsenhuber et al., 2020) and have also led to the development of scaling laws to estimate the extent of atmospheric erosion in such events (e.g., Shuvalov, 2009; Kegerreis et al., 2020b; Denman et al., 2022). These studies commonly suggest that only partial atmospheric erosion occurs in typical giant impacts. For instance, the atmospheric mass loss prescription from Kegerreis et al. (2020b) estimates that only $10\%$ of Earth’s primordial atmosphere could have been lost from the immediate effects of the canonical Moon forming impact.
@@
Due to the prohibitive computational cost associated with these numerical simulations, the primary focus of these studies has typically been on the impact and its immediate aftermath, typically covering only a few days. However, these impacts generally eject a fraction of mass from the colliding bodies, some of which persists within the system long after the primary impact (e.g., Jackson & Wyatt, 2012; Jackson et al., 2014; Genda et al., 2015; Watt et al., 2024). Depending on the specifics of the impact, some fraction of this ejected mass would be vaporised (e.g., Lock & Stewart, 2017; Carter et al., 2020; Davies et al., 2020; Watt et al., 2021; Nakajima et al., 2022; Caracas & Stewart, 2023). As the vaporised ejecta expands, it cools and condenses, producing dust at detectable levels, which are then rapidly depleted to non-detectable levels through collisional evolution and subsequently removed from the system via radiation forces (e.g., Wyatt & Jackson, 2016; Watt et al., 2021). The non-vaporised debris survives much longer and will gradually erode through mutual collisions and other physical processes, including gravitational perturbations from the surviving planet(s) (e.g., Jackson & Wyatt, 2012; Watt et al., 2024). Over time, these objects will be progressively ground down to micrometre-sized dust through collisional evolution and eventually removed by radiation pressure or Poynting-Robertson drag (e.g., Wyatt, 2008; Wyatt et al., 2011). This theoretical understanding is supported by observations of extrasolar systems with large quantities of warm dust in the terrestrial region, inferred to have been recently created in collision between the planetary embryos (e.g., Lisse et al., 2009; Weinberger et al., 2011; Genda et al., 2015; Leinhardt et al., 2015; Su et al., 2019; Moór et al., 2021; Schneiderman et al., 2021; Kenworthy et al., 2023; Watt et al., 2024; Su et al., 2025). Importantly, during this long process, a fraction of the debris will re-impact the planet and be re-accreted into the remaining planet(s) (Jackson & Wyatt, 2012; Wyatt et al., 2017), potentially causing further atmospheric erosion.
@@ 2 Methods
In this study, we adopt a simplified approach to investigate the long-term consequences of giant impacts. Depending on the kinematics, giant impacts can result in a range of different outcomes, from efficient accretion and hit-and-run collisions to catastrophic disruptions (e.g., Leinhardt & Stewart, 2012; Emsenhuber et al., 2020). In all classes of impact we consider, a fraction of the parent body mass becomes unbound and is ejected. This unbound mass is generally a mixture of vapour and solid material [^1] By solid material, we refer to all the non-vaporised condensed matter, which may eventually cool to become solid debris. with varying proportions, depending on the specifics of the impact (e.g., Ćuk & Stewart, 2012; Lock & Stewart, 2017; Carter et al., 2020; Davies et al., 2020; Watt et al., 2021; Nakajima et al., 2022). However, in this study, our focus is not on the detailed outcomes of individual giant impact events. Instead, we attempt to parametrize the outcome in a simplified manner that is broadly applicable across a range of giant impacts. Hence, we adopt a simplified model in which a giant impact results in the ejection of a fraction $f_{\rm{esc}}$} of the surviving planet’s mass $m_{\rm{p}}$} as unbound solid material and a fraction $f_{\rm{esc,v}}$} as vapor. The unbound ejecta would expand out from the point of collision and shear out to eventually form an axisymmetric debris disk around the host star encompassing the planet and the collision point (e.g., Jackson & Wyatt, 2012; Jackson et al., 2014; Genda et al., 2015; Watt et al., 2024). The debris are assumed to have some velocity dispersion in addition to the Keplerian velocity of the planet.
@@ 2.1 Evolution of Debris from Giant Impacts & Re-accretion
where, $M_{\rm{*}}$} is the stellar mass, $a_{\rm{p}}$} is the semi-major axis, and $\rho_{\rm{p}}$} is the bulk density of the planet. Throughout the study, we have assumed that the debris has a density of ${\rho_{\rm{imp}}}=2.7\ {{\rm g\ cm^{-3}}}$, and a dispersal threshold ${Q_{\rm{D}}^{*}}=10^{5}\ {{\rm J\ kg^{-1}}}$. The size of the largest objects in the debris ($D_{\rm{max}}$}) from a giant impact is highly uncertain. The vaporised component of the ejecta would expand and cool to form condensates with sizes ($D_{\rm{vc}}$}) in the millimetre to centimetre range (e.g., Melosh & Vickery, 1991; Johnson & Melosh, 2014), while the non-vaporised material may reach up to several hundred kilometres. For instance, Genda et al. (2015) found that the largest debris fragments [^3] Here, debris fragments refer to the unbound ejecta, not to the remnant post-impact protoplanet(s). could be as massive as half the lunar mass, corresponding to a diameter of $\sim 1380\ {\rm{km}}$. More recently, Watt et al. (2024) found that the debris fragments are generally larger than $\sim 100\ {\rm{km}}$ in diameter. This suggests that the size distribution of the ejecta should consist of two primary components, the smaller vapour condensates and the larger non-vaporised material, with their relative proportions and the sizes of the largest non-vaporised bodies depending on the specific impact conditions. Given the substantial differences in their sizes, we expect each component to set up a collisional cascade and to evolve essentially independently of the other. The small vapour condensates would have short collisional lifetimes, and are therefore rapidly depleted from the system, typically within a small number of orbits (see Figure 5 of Wyatt & Jackson, 2016). Due to their small size, these condensates are ineffective in atmospheric erosion (e.g., Schlichting et al., 2015) and thus would have a negligible influence on the post-impact planet(s) over their short lifetimes. In contrast, the larger non-vaporised debris would have much longer collisional lifetimes (e.g., $\tau_{c}\sim 6$ Myr for ${D_{\rm{max}}}=100\ {\rm{km}}$, considering an Earth-like progenitor at $1\ {\rm{au}}$ and ${M_{\rm{d}}}=0.01\ {M_{\earth}}$), increasing the likelihood of re-accretion and potentially influencing the planet’s subsequent evolution. Therefore, we exclude the vaporised component from our analysis and, for the remainder of this study, focus exclusively on the non-vaporised debris, adopting ${M_{\rm{d}}}={f_{\rm{esc}}}{m_{\rm{p}}}$.
@@ 3 Results
We first consider the possible implications of debris re-accretion on Earth following a giant impact such as that hypothesised to have formed the Moon. This extensively studied event, observationally constrained by the Earth-Moon configuration (e.g., Cameron & Ward, 1976; Canup & Asphaug, 2001; Timpe et al., 2023; Zhou et al., 2024), is a natural reference point for our study. We note, however, that the exact nature of the impact scenario that led to the formation of the Moon remains uncertain and debated. It is not the purpose of our study to address these, but to adopt one representative scenario as a reference. While we use the giant impact hypothesis of the Moon formation as a reference point, the focus of our work is not restricted to satellite-forming impacts. Instead, we aim to explore the broader role of typically overlooked debris in causing atmospheric erosion through secondary impacts, regardless of whether the primary impact forms a satellite or not. To this end, we subsequently expand our analysis to explore a broader range of parameter space.
@@ 3.1.3 Effect of Initial Atmospheric Properties
In the previous sections, we assumed that Earth, after the giant impact, had an atmosphere similar to that of the present-day Earth. However, the initial atmospheric and surface conditions can have a significant effect on the subsequent atmospheric erosion (e.g. Lock & Stewart, 2024). In this section, we explore the long-term atmospheric erosion following the primary impact, focusing on the effects of initial atmospheric mass and mean molecular weight on this process.
@@
The mean molecular weight of the atmosphere, $\mu$, directly affects the atmospheric scale height $H$ ($H\propto 1/\mu$; Equation 9), while the atmospheric mass primarily influences the base density, $\rho_{\rm{b}}$}. For a given planet and impact velocity, the values of $H$ and $\rho_{\rm{b}}$} determine the size range of the impactors that most efficiently erode the atmosphere. In the atmospheric erosion prescription of Shuvalov (2009), this relationship is encapsulated in the definition of the erosional efficiency factor, $\xi$ (section 2.2), which determines the extent of atmospheric erosion (Equation 11). For a given impact velocity and impactor mass, atmospheric mass loss peaks at $\xi\sim 350$ and falls for both higher and lower values of $\xi$, declining to less than $10\%$ of its peak value beyond the range $10\lesssim\xi\lesssim 10^{5}$ (see Figure 6 of Shuvalov (2009)). From section 2.2, we find that $D^{3}\propto\xi{m_{\rm{atm}}}/\mu^{2}$. So, as $m_{\rm{atm}}$} increases, the efficient erosion regime ($10\lesssim\xi\lesssim 10^{5}$) shifts towards the larger impactor sizes, where we have more available mass in the debris, based on the adopted debris size distribution (section 2.1). If the efficient erosion regime remains below the maximum debris size ($D_{\rm{max}}$}), the atmospheric erosion per unit accreted mass increases (Equation 14). However, a continued increase in $m_{\rm{atm}}$} may shift the efficient erosion regime to even larger impactor sizes exceeding $D_{\rm{max}}$}, leading to a decline in atmospheric erosion. This transition occurs at lower $m_{\rm{atm}}$} for lower values of $\mu$ (e.g., at ${m_{\rm{atm}}}\simeq 0.7\ {m_{\rm{atm\earth}}}$ for $\mu=2.35$, while at ${m_{\rm{atm}}}\simeq 100\ {m_{\rm{atm\earth}}}$ for $\mu=29$) due to the $\mu$ dependence of $\xi$.